OpenStudy (anonymous):

If f'(x)= cos x, and f(pi/2)=-2, find f(x).

jimthompson5910 (jim_thompson5910):

can you think of what goes in the blank space (inside the square brackets)? \[\Large \frac{d}{dx}\left[ \ \ \ \ \ \right] = cos(x) \]

OpenStudy (solomonzelman):

take the indefinite integral of cos(x), and then plug in the point into your anti-derivative function to find what your C is.

OpenStudy (bradely):

f'(x) = cos x f(x) = sin x+C f(pi/2) =-2 sin pi/2+C =-2 C = -3 f(x) = sin x-3 Source: http://www.mathskey.com/question2answer/